### v1

# Multi-view Graph Structural Representation Learning via Graph Coarsening

Abstract
Graph Transformers (GTs) have made remarkable achievements in graph-level tasks. However, most existing works regard graph structures as a form of guidance or bias for enhancing node representations, which focuses on node-central perspectives and lacks explicit representations of edges and structures. One natural question is, can we treat graph structures node-like as a whole to learn high-level features? Through experimental analysis, we explore the feasibility of this assumption. Based on our findings, we propose a novel multi-view graph structural representation learning model via graph coarsening (MSLgo) on GT architecture for graph classification. Specifically, we build three unique views, original, coarsening, and conversion, to learn a thorough structural representation. We compress loops and cliques via hierarchical heuristic graph coarsening and restrict them with well-designed constraints, which builds the coarsening view to learn high-level interactions between structures. We also introduce line graphs for edge embeddings and switch to edge-central perspective to construct the conversion view. Experiments on six real-world datasets demonstrate the improvements of MSLgo over 14 baselines from various architectures.
1 Introduction
Graph Neural Networks (GNNs) have become a significant approach to Graph Representation Learning (GRL) recently, achieving remarkable performance on various node-level (such as node classification Chen et al. (2023a) and link prediction Chamberlain et al. (2023)) and graph-level (such as graph classification Niepert et al. (2016); Xinyi and Chen (2019); Yao et al. (2023)) tasks. GNNs model the graph structure implicitly, leveraging graph topology information to aggregate neighborhood node features. However, Xu et al. Xu et al. (2019) have proved that the aggregation mechanism of GNNs has local limitations, which has encouraged more researchers to move from GNNs to GTs Liu et al. (2023a); Chen et al. (2023b); Wu et al. (2023). GTs utilize position encoding to represent the graph topology globally, jumping out of the neighborhood restriction and interacting between distant node pairs. Although GTs can learn long-distance structural information, most GNNs and GTs use graph structures as guidance Huo et al. (2023) or bias Ying et al. (2021) to obtain better node representations rather than directly representing them.
A-BA-CA-DB-CB-DC-D000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Graph PairsEuclidean Distanceloopdiffgraph0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Cosine Similarityloopdiffgraph
ABCD000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Graph PairsEuclidean Distanceloopdiffgraph0.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.91111Cosine Similarityloopdiffgraph
Figure 1: Pre-experiment on a toy molecule graph set. (a) shows four graphs with different features, all containing a benzene ring. (b) illustrates the pairwise comparison between four graphs. (c) presents the comparison between the graph and its coarsening view. We leverage two measurements of Euclidean Distance (bar plots, lower for closer) and Cosine Similarity (line plots, higher for closer) to reveal the latent relations between graphs. Best viewed in color.
Structural information is crucial for graph-level tasks, especially Graph Structural Learning (GSL) Zhang et al. (2023a); Zou et al. (2023); Zhang et al. (2023c). GSL aims to learn the optimized graph structure and representation from the original graph, further improving the model performance on downstream tasks (e.g. graph classification). Kernel-based methods Shervashidze et al. (2009, 2011) lead the way of GSL in the early stage, which selects representative graph structures carefully to test the isomorphism and update weights in classifiers. However, this category of methods suffers due to computational bottlenecks of the pair-wise similarity Wu et al. (2021); Yang et al. (2023). With the development of deep learning, researchers have proposed lots of new graph learning paradigms such as Graph Contrastive Learning (GCL) Sun et al. (2020); Ma et al. (2023); Liu et al. (2023b). Using various data augmentation strategies like node dropping Ma et al. (2023), edge perturbation Zhang et al. (2023b), attribute masking Yuan et al. (2023), and subgraph sampling Liu et al. (2023b), GCL ensures the semantic matching between views and enhances the robustness of the model against multifaceted noises. Though GCL has made considerable achievements, views obtained through the strategies above keep the original structures, which restricts model capabilities since limited usage of high-level structural information.
To fully utilize topological information and capture high-level structures, one natural question arises as to whether we can treat graph structures node-like as a whole for graph-level tasks. Take the molecule graph as an example, specific structures such as molecular fragments and functional groups contain rich semantics. Random perturbation of these structures will produce additional structural information, which has been proved to be invalid nevertheless Zhu et al. (2023). Some researchers have focused on this problem and defined the overall treatment idea as Graph Compression Liu et al. (2022), Pooling Wei et al. (2021); Baek et al. (2021); Lv et al. (2023), and Coarsening Zhang et al. (2023b). Inspired by Convolutional Neural Networks (CNNs) to compress the set of vectors into a compact representation Wu et al. (2022a), this series of works aims to address the flat architecture of GNNs, emphasizing the hierarchical structures of graphs Ying et al. (2018).
We further explore the feasibility of treating graph structures node-like as a whole through a set of pre-experiments. Figure 1 shows the results. The pre-experiments follow the plain message-passing mechanism and summation readout function. We train the pre-experimental model on four toy graphs and obtain representations of the whole graph, benzene ring structure, and main distinguishing atom sets. We measure the spatial distance and angular separation between representations by Euclidean Distance and Cosine Similarity, respectively. As observed in Figure 1b, the gap between loop is less, w.r.t. graph and diff. It highlights that the prime factor causing the differences lies behind the main distinguish atom sets rather than the benzene ring. Thus, we compress the benzene ring into one node with the same settings and obtain the coarsening view of each graph. Figure 1c shows the distance between the original graph and its coarsening view. Pairs such as A and C, whose main distinguish atom sets are arranged adjacent, leave a relatively closer cosine similarity, while B and D show a significant difference. It indicates that simple coarsening will lose part of structural information. In summary, pre-experiments tell us two enlightenments: (1) some structures contribute relatively less to distinguish graphs, which can further turn into a coarsening view to magnify the high-level structural information; (2) after treating structures node-like as a whole, some of the structural information may be traceless, calling for additional consideration for the relative position of neighborhoods.
Therefore, based on the above discovery, we focus on loops and cliques and introduce two views: a coarsening view and a conversion view to emphasize individual components of the graph. The former coarsens graph structures into one node based on clustering to learn high-level structural information. The latter transforms nodes and edges from the node-central perspective to edge-central, which highlights relative position and is complementary to the loss of coarsening. Specifically, we construct the graph coarsening view through a heuristic algorithm restricted by well-designed constraints and build the line graph conversion view to augment relative position information. Together with the original graph view, we train separate GT for each view for graph encoding. Finally, we concatenate multi-view representation vectors as the final embedding of an entire graph.
The contributions of this paper are as follows:
(1) We design a novel multi-view graph structural representation learning model via graph coarsening (MSLgo), which leverages three views to learn more comprehensive structural information for downstream graph classification tasks.
(2) We propose a hierarchical heuristic algorithm to compress the loops and cliques and construct the graph coarsening view, which captures high-level topological structures.
(3) We introduce the line graph conversion view, which retains relative position information between neighbor nodes.
(4) We verify the performance of MSLgo on 6 datasets from 2 categories, compared with 14 baselines from 5 categories. MSLgo achieves better results than SOTAs.
Figure 2: Framework of our proposed model MSLgo (top center). Blocks below are the loop structure coarsening view, clique structure coarsening view, and line graph conversion view from left to right. Best viewed in color.
2 Related Work
Graph Pooling.
Graph Pooling Wei et al. (2021); Baek et al. (2021); Lv et al. (2023) is one of the pivotal techniques for graph representation learning, aiming to compress the input graph into a smaller one. Due to the hierarchy characteristics of graphs, hierarchical pooling methods have become mainstream nowadays. Methods fall into two categories depending on whether breaking nodes. The first category conducts new nodes via graph convolution. The pioneer in hierarchical pooling, DIFFPOOL Ying et al. (2018), aggregates nodes layer by layer to learn high-level representations. Inspired by the Computer Vision field, CGKS Zhang et al. (2023b) constructs the pooling pyramid under the contrastive learning framework, extending the Laplacian Eignmaps with negative sampling. MPool Islam et al. (2023) takes advantage of motifs to capture high-level structures based on motif adjacency. The second category condenses graphs by sampling from old nodes. SAGPool Lee et al. (2019) masks nodes based on top-rank selection to obtain the attention mask subgraphs. CoGSL Liu et al. (2022) extracts views by perturbation, reducing the mutual information for compression. MVPool Zhang et al. (2023c) reranks nodes across different views with the collaboration of attention mechanism, then selects particular node sets to preserve underlying topological.
Graph Transformers.
Graph Transformers (GTs) Chen et al. (2023b); Wu et al. (2023) alleviate the problems of over-smoothing and local limitations of GNNs, which has attracted the attention of researchers. The self-attention mechanism learns long-distance interactions between every node pair and shows tremendous potential for various scenarios. Graphormer Ying et al. (2021) integrates edges and structures into Transformers as biases, outperforming in graph-level prediction tasks. U2GNN Nguyen et al. (2022) alternates aggregation function leveraging Transformer architecture. Exphormer Shirzad et al. (2023) builds an expansive GT leveraging a sparse attention mechanism with virtual node generation and graph expansion. In addition, due to the quadratic computational constraint $O(N^{2})$ of self-attention, several works Wu et al. (2022b); Liu et al. (2023a) have been proposed to focus on a scalable and efficient Transformer. The former puts forward a new propagation strategy adapting arbitrary nodes in scale. The latter combines pooling blocks before multi-head attention to shrink the size of fully connected layers. To summarize, we design our model based on the capabilities of GTs, which can effectively aggregate features of distant nodes using attention mechanisms and overcome the limitations of local neighborhoods.
3 Preliminaries
3.1 Notations
Graphs.
Given a graph $G=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denote the set of nodes and edges, respectively. We leverage $A\in\{0,1\}^{N\times N}$ to indicate the adjacency matrix, and $A[i,j]=1$ when there is an edge between node $v_{i}$ and $v_{j}$.
Problem Definition.
For supervised graph representation learning, given a set of graphs $\mathcal{G}=\{G_{1},\cdots,G_{M}\}$ and their labels $\mathcal{Y}=\{y_{1},\cdots,y_{M}\}$, our goal is to learn a representation vector $H_{i}$ for each graph, which can be used in downstream classification tasks to predict the label correctly.
3.2 Universal Graph Transformer (U2GNN)
U2GNN Nguyen et al. (2022) is a GNN-based model that follows the essential aggregation and readout pooling functions. Xu et al. Xu et al. (2019) claim that a well-designed aggregation function can further improve the performance. Thus, U2GNN replaces the plain strategy which GNN uses with the following aggregation function:
$h_{t,v}^{(k)}=\mathbf{AGG}_{\mathsf{att}}(h_{t-1,v}^{(k)}),$ (1)
where $h_{t,v}^{(k)}$ denotes the representation vector of node $v$ in step $t$ of layer $k$, which aggregates from step $t-1$. It provides a powerful method based on the transformer self-attention architecture to learn graph representation. We take U2GNN as the backbone and cover the details later in Section 4.4.
4 Methodology
4.1 Overview
Figure 2 shows the framework diagram of our proposed model MSLgo, which mainly contains three parts: multi-view construction of graphs, graph encoders, and classifier (omitted in the figure). Given an input graph, MSLgo first constructs three views: the original graph view, graph coarsening view, and line graph conversion view, via designed methods. Subsequently, MSLgo trains a GT for each view separately to obtain graph-level representation, which is directly concatenated and fed into the downstream classifier.
4.2 Loop and Clique Coarsening
From pre-experiments, we know that treating graph structures node-like as a whole via graph coarsening is possible. Existing methods mainly pay attention to the hierarchical structure of graphs: some cluster nodes based on graph convolution to do graph coarsening Ying et al. (2018), and some implement node condensing by iterative sampling Lee et al. (2019). Aggregation of first-order adjacent nodes has intuitive interpretability, which means connections between neighborhoods, while a deeper coarsening shows a poor explanation. Moreover, in high-level coarsening, the main distinguishing node sets disappear into clusters, which leads to an ambiguity for topology. Thus, MSLgo achieves graph coarsening of loops and cliques with shallow coarse-grained clustering restricted by two hard constraints, which magnifies high-level structural information while preserving the characteristic nodes in graphs to the greatest extent.
We first give formal definitions of loops and cliques.
Definition 1 (Loop).
Given an undirected graph $G$, a path $G_{l}=\{v_{1}v_{2}\cdots v_{n}v_{n+1}\}$ is a loop if there is no repetitive node in $\{v_{1},\cdots,v_{n},v_{n+1}\}$ except $v_{1}=v_{n+1}$ and $(v_{k},v_{k+1})\in\mathcal{E}_{G}$ for $k=1,2,\cdots,n$. $|\mathcal{V}_{G_{l}}|$ is the length of the loop.
Definition 2 (Clique).
Given an undirected graph $G$, a subgraph $G_{c}$ is a clique if $\forall v_{i},v_{j}\in\mathcal{V}_{G_{c}},(v_{i},v_{j})\in\mathcal{E}_{G}(i\neq j)$. $G_{c}$ is a k-clique if $|\mathcal{V}_{G_{c}}|=k$.
Algorithm 1 Loop and Clique Coarsening (LCC)
Input: the input graph $G_{i}$, the max loop length $\delta$, the max hierarchy depth $\sigma$;
Output: the coarsening view $G_{co}$
1:  $G_{co}\xleftarrow{}G_{i}$;
2:  $vis\xleftarrow{}\emptyset,d\xleftarrow{}0$;
3:  find node set $cur$ in $G_{co}$ with the highest degree;
4:  while not all nodes in vis or $d<\sigma$ do
5:     fetch $G_{c}$ not visited containing nodes $c\subset cur$;
6:     $G_{co}\xleftarrow{update}G_{c}$;
7:     $vis\xleftarrow{}vis\cup G_{c}$;
8:     $cur\xleftarrow{}cur-c$;
9:     if $cur=\emptyset$ then
10:        find new $cur$ not visited in next depth $d+1$;
11:        $d\xleftarrow{}d+1$;
12:     end if
13:  end while
14:  if less or no cliques found then
15:     $G_{lt}\xleftarrow{}\emptyset$;
16:     for all $v_{i}\in G_{co}$ do
17:        find all loops $L_{i}$ in $\delta$ length beginning with $v_{i}$;
18:        $G_{lt}\xleftarrow{update}L_{i}$;
19:     end for
20:     $G_{co}\xleftarrow{}G_{lt}$;
21:  end if
22:  return $G_{co}$
Coarsening requires counting all loops and cliques in the graph. However, according to the definitions above, loops are contained within cliques. Moreover, since the Maximum Clique Problem (MCP) is NP-hard Bomze et al. (1999), finding all cliques is also NP-hard. Algorithms require an approximation or a shortcut due to enormous search space.
Algorithm 1 describes the Loop and Clique Coarsening (LCC) process. Based on Tarjan Trajan (1974) and Clique Percolation (CP) Palla et al. (2005), LCC heuristically iterates the graph hierarchy using loop length and hierarchy depth as constraints for pruning. LCC finds cliques under depths less than $\sigma$ (lines 4-8) first, counting loops only when finding fewer or no cliques (lines 9-16). When updating graphs (lines 6 and 13), LCC reconstructs graphs to build the coarsening view. Mathematically, a new node $\bar{v}$ sums all node features in the structure (loop or clique) $S$ to represent it as follows:
$h_{\bar{v}}=\sum_{v_{i}\in S}h_{v_{i}},$ (2)
where $h_{*}$ denotes the feature of *.
Example 1.
We take Figure 2 as an example. Given an input graph $G_{i}$ (top left), we will get a reconstructed graph (top right) after coarsening loops and cliques.
As for the limitation of search complicity, we empirically set two constraints of loop length and hierarchy depth to prune.
Loop Length Constraint.
Not all loops are suitable for coarsening. Long loops may reveal two chains or sequences interacting (e.g., backbone in proteins Wang et al. (2023)). The semantics will change if coarsening happens. MSLgo sets a range $\delta$ for the maximum detection length (line 13). We default it to 6 to cover usual cases, such as squares and benzene rings.
Example 2.
The Loop Length Constraint in Figure 2 (bottom left) illustrates the case of $\delta=4,5$. Only loops with length $3<l\leq\delta$ are counted. It skips loops with length over $\delta$ and pushes triangles (3-clique) to Clique Coarsening.
Hierarchy Depth Constraint.
The NP-hard nature of finding all cliques determines the impossibility of exhaustive search, especially for large graphs. MSLgo takes advantage of the hierarchical characteristic of graphs, taking nodes with the highest degree (lines 3 and 7) hop by hop to find cliques. We set a distance $\sigma$ to control the depth of the recursion.
Example 3.
The Hierarchy Depth Constraint in Figure 2 (bottom center) shows choices of the yellow center node after graph coarsening with $\sigma=1$. It compresses the red clique and blue node in one hop while preserving the yellow one. Nodes at $d>\sigma$ stay still.
4.3 Line Graph Conversion
This section tells why to build the Line Graph Conversion (LGC) view. Line graph is a usual method for edge representations in node-level tasks such as POI recommendation Chanpuriya et al. (2023). Edge augmentation is a considerable view to make a thorough representation integrating various architectures and scenarios Mo and Yamana (2023); Liu et al. (2023a). Therefore, we build LGC to attach an edge-central perspective in graph-level tasks. Firstly, the definition of line graphs is as follows.
Definition 3 (Line Graph).
Given an undirected graph $G=\{\mathcal{V},\mathcal{E}\}$, $L(G)=\{L(\mathcal{E}),\mathcal{E}_{L(G)}\}$ is the line graph of $G$, where $L(\mathcal{E})$ and $\mathcal{E}_{L(G)}$ denote the node set and edge set of the line graph, and $L$ is a mapping function from the edge set of the original graph to the node set of the line graph, satisfying $\forall e_{ij}\in\mathcal{E},L(e_{ij})\in L(\mathcal{E})$, and $\forall e_{ij},e_{ik}\in\mathcal{E},(L(e_{ij}),L(e_{ik}))\in\mathcal{E}_{L(G)}(j\neq k)$.
In plain words, a line graph turns the edges in the original graph into new vertices and the endpoints of two edges in the original graph into new edges. This explicit transformation can reduce the difficulty of relative position modeling, which is implicit in the node-central view of the original graph. We can build LGC through existing tools Hagberg et al. (2008) conveniently. For those graphs with no edge attributes, we sum features of endpoints for consistency following Equation 2. Moreover, we substitute dense graph inputs and large-scale graphs with multi-coarsening ones for lower time and space consumption.
There exists relative position information loss once we treat graph structures node-like as in previous section. LGC is a supplementary of the coarsening view. We explain this from two aspects: different position information of neighborhood nodes and some hard-coarsening examples.
Different Position Information.
From pre-experiments, we know that relative position suffers if we condense graph structures into one node. Though plain coarsening strategies decrease the graph scale, it makes positions vague, such as between Graphs A and B in Figure 1a. In essence, position information is the relative position among nodes connected by node-edge sequence. LGC models edge-central positional relationships, describing structural information by adding a new view to obtain a more comprehensive representation.
Example 4.
The Different Position Information in Figure 2 (bottom right) shows two similar graphs with a slight difference (the relative position of the edges is bold in red and blue). Coarsening the square will eliminate the difference, while LGC can enhance the diversity.
Hard Coarsening Examples.
MSLgo only focuses on loops and cliques. However, not all graphs contain either structures, such as long chains. A simple copy view contributes less if we leave these hard-coarsening examples alone. LGC transforms these graphs into a fittable form that is easy for postprocessing while preserving the original structural information.
Example 5.
The Hard Coarsening Examples in Figure 2 (bottom right) list two cases: claw-like structures (blue) and 4-node chains (red). Both are hard to coarsen since neither loop nor clique. After LGC, the former turns into a triangle, and the latter decreases the chain length. Both of the structures become easier for further processing.
4.4 Our Model: MSLgo
After acquiring three views, we can build our model MSLgo based on U2GNN Nguyen et al. (2022). As shown in Figure 2, MSLgo takes $G_{i}$ as input and obtains the coarsening view $G_{co}$ and line graph conversion view $L(G_{i})$ with designed algorithms. For $G_{co}$, MSLgo calculates node embeddings $h$ according to Equation 1 and uses the summation readout function to obtain graph embedding $H_{co}$. Specifically, it is
$H_{co}=\sum_{v\in G_{co}}[h_{0,v}^{1};\cdots;h_{0,v}^{K}],$ (3)
where $h_{0,v}^{k}$ denotes the initial embedding of node $v$ in layer $k$. For each layer $k$, MSLgo iterates $T$ steps aggregating $N$ sampling neighbors and passes $h_{T,v}^{k}$ to the next layer $k+1$.
We use the same operation for the other two views, $G_{o}$ and $L(G_{i})$, and achieve corresponding embeddings, $H_{o}$ and $H_{lG}$. We apply concatenation across views to obtain the final embedding $H_{i}$ for the input graph $G_{i}$ as
$H_{i}=[H_{o};H_{co};H_{lG}].$ (4)
Apart from a single fully connected (FC) layer, we vary the classifier with a 2-layer MLP,
$\hat{y}_{i}=W_{2}(W_{1}H_{i}+b_{1})+b_{2},$ (5)
where $W_{1}$ and $W_{2}$ are the weight matrices, and $b_{1}$ and $b_{2}$ are the bias parameters, adapting to the embedding dimension increase. The loss function is cross-entropy as follows:
$loss=-\sum_{i=1}^{M}y_{i}\log(\sigma(\hat{y}_{i})),$ (6)
where $\sigma$ denotes the softmax function. For further implementation details, we suggest to refer to  Nguyen et al. (2022).
5 Experiments
5.1 Experimental Settings
Datasets.
We evaluate our approach on six widely used datasets from TUDataset Morris et al. (2020), including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and NCI1 (N1). Table 1 shows the details of the datasets.
Dataset # Graphs # Classes Avg.N Avg.E
CO 5000 3 74.49 2457.78
IB 1000 2 19.77 96.53
IM 1500 3 13.00 65.94
DD 1178 2 284.32 715.66
N1 4110 2 29.87 32.30
PTC 344 2 14.29 14.69
Table 1: The statistics of datasets from TUDataset. Avg.N and Avg.E denote the average number of nodes and edges.
Method CO IB IM DD N1 PTC
GK 72.84$\pm$0.28 65.87$\pm$0.98 43.89$\pm$0.38 78.45$\pm$0.26 62.30$\pm$0.30 57.26$\pm$1.41
WL 79.02$\pm$1.77 73.40$\pm$4.63 49.33$\pm$4.75 79.78$\pm$0.36 82.19$\pm$0.18 57.97$\pm$0.49
GAT 75.80$\pm$1.60 70.50$\pm$2.30 47.80$\pm$3.10 - 75.20$\pm$3.30 66.70$\pm$5.10
GMT 80.74$\pm$0.54 73.48$\pm$0.76 50.66$\pm$0.82 78.72$\pm$0.59 - -
MSGNN - 79.21$\pm$1.07 57.38$\pm$1.09 87.29$\pm$2.17 76.89$\pm$3.23 -
PSCN 72.60$\pm$2.15 71.00$\pm$2.29 45.23$\pm$2.84 77.12$\pm$2.41 76.34$\pm$1.68 62.29$\pm$5.68
GraphSAGE 79.70$\pm$1.70 72.40$\pm$3.60 49.90$\pm$5.00 65.80$\pm$4.90 - 63.90$\pm$7.70
DGCNN 73.76$\pm$0.49 70.03$\pm$0.86 47.83$\pm$0.85 79.37$\pm$0.94 74.44$\pm$0.47 58.59$\pm$2.47
CapsGNN 79.62$\pm$0.91 73.10$\pm$4.83 50.27$\pm$2.65 75.38$\pm$4.17 78.35$\pm$1.55 66.00$\pm$1.80
SEK-GIN - 76.90$\pm$3.80 53.80$\pm$3.20 - - 66.86$\pm$7.20
InfoGraph 70.65$\pm$1.13 73.03$\pm$0.87 49.69$\pm$0.53 72.85$\pm$1.78 73.80$\pm$0.70 61.65$\pm$1.43
GraphACL 74.72$\pm$0.55 74.53$\pm$0.39 51.65$\pm$0.34 79.34$\pm$0.42 - -
CGKS 76.80$\pm$0.10 - - - 79.10$\pm$0.20 63.50$\pm$1.30
U2GNN 77.84$\pm$1.48 77.04$\pm$3.45 53.60$\pm$3.53 80.23$\pm$1.48 69.81$\pm$2.36 69.63$\pm$3.60
MSLgo 97.60$\pm$0.81 88.40$\pm$2.73 63.60$\pm$2.83 93.21$\pm$3.26 79.34$\pm$1.88 71.53$\pm$6.59
Table 2: The main results of the graph classification task on six datasets (mean accuracy (%) and standard deviation). The best scores are in bold. ’-’ indicates that results are unavailable in the original or published papers.
Baselines.
To evaluate the effectiveness of our model, we select 14 related works from 5 categories for comparison. (I) 2 kernel-based methods: GK Shervashidze et al. (2009) and WL Shervashidze et al. (2011); (II) 3 attention-based and pooling methods: GAT Velickovic et al. (2018), GMT Baek et al. (2021), and MSGNN Lv et al. (2023); (III) 5 GNN-based methods: PSCN Niepert et al. (2016), GraphSAGE Hamilton et al. (2017), DGCNN Zhang et al. (2018), CapsGNN Xinyi and Chen (2019), and SEK-GIN Yao et al. (2023); (IV) 3 contrastive learning methods: InfoGraph Sun et al. (2020), GraphACL Luo et al. (2023), and CGKS Zhang et al. (2023b); and (V) 1 backbone: U2GNN Nguyen et al. (2022) with Transformer style.
Implementation Details.
We follow the work Nguyen et al. (2022) to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. In detail, we set the batch size to 4 and the dropout to 0.5. We pick the hierarchy depth in $\{1,2,3\}$, the number of Transformer Encoder layers in $\{1,2,3,4\}$, the number of sampling neighbors in $\{4,8,16\}$, the hidden size in $\{128,256,512,1024\}$, and the initial learning rate in $\{5e^{-4},1e^{-4},5e^{-4},1e^{-3}\}$. We utilize Adam Kingma and Ba (2015) as the optimizer. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM GPU.
(a) Boxplot and Loss
(b) 10 folds in last epoch
Figure 3: Robustness studies on the CO dataset.
5.2 Results and Analysis
Main Results.
Table 2 shows the main results of the graph classification task. MSLgo outperforms all other baseline models on all datasets except N1, where we achieve a suboptimal performance. In detail, the boost of MSLgo is higher in social networks than in bioinformatics. The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. The improvement of PTC is only 1.90%, which concludes that long chain structures make the compression harder, where fewer loops and cliques appear.
Robustness Studies.
We also investigate the robustness of MSLgo. Figure 3 shows the boxplot of the performance for 10-fold and the average loss under each epoch of training on the CO dataset. It indicates that the performance stabilizes as the training progress. On the one hand, for every 10-fold, the standard deviation of the performance gradually decreases, especially for the last 5 epochs, where the accuracy stays above 95%. On the other hand, there are outliers (mainly lower performance) in some epochs. This phenomenon only occurs in the first half of training and disappears after Epoch 26, which explains that the learning ability of MSLgo becomes stable and no longer fluctuates.
Method CO IB IM
MSLgo 97.60$\pm$0.81 88.40$\pm$2.73 63.60$\pm$2.83
w/o LG 97.46$\pm$0.57 87.60$\pm$3.20 60.73$\pm$3.68
w/o GC 89.66$\pm$1.56 82.60$\pm$3.10 59.40$\pm$3.99
U2GNN 77.84$\pm$1.48 77.04$\pm$3.45 53.60$\pm$3.53
Method DD N1 PTC
MSLgo 93.21$\pm$3.26 79.34$\pm$1.88 71.53$\pm$6.59
w/o LG 91.43$\pm$2.77 78.73$\pm$1.28 71.25$\pm$5.39
w/o GC 83.78$\pm$4.63 72.38$\pm$2.14 70.67$\pm$5.21
U2GNN 80.23$\pm$1.48 69.81$\pm$2.36 69.63$\pm$3.60
Table 3: Ablation studies on all six datasets, social (upper) and bioinformatics (below).
5.3 Ablation Study
[…]
5.4 Hyper-parameters Study
In this section, we explore the hyper-parameters of our model. We conduct experiments with different settings for 4 hyper-parameters: $K$, $T$, $N$, and hidden size on 3 datasets: IB, IM, and PTC. Figure 4 shows the results of hyper-parameters studies.
[…]
(a) $K$: hierarchy depth
[…]
(b) $T$: the number of layers
[…]
(c) $N$: sampling neighbors
777788889999101010100.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9Hidden Size ($2^{x}$)AccuracyIBIMPTC
(d) Hidden Size
Figure 4: Hyper-parameters studies. Best viewed in color.
Hierarchy Depth $K$ and Layers $T$.
As shown in Figure 4a and Figure 4b, MSLgo has different choices in the hierarchy depth $K$ and the number of Transformer Encoder layers $T$. In short, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy.
Sampling Neighbors $N$ and Hidden Size.
Figure 4c and Figure 4d show the performance of MSLgo when sampling different numbers of neighbors and selecting different hidden layer sizes, respectively. We can see from the results that the larger the number of sampling neighbors and hidden size, the better the model performance. It indicates that the increase in the sampling and hidden size within a particular range can promote the model to learn high-level structures thoroughly and clearly.
5.5 Case Study
Refer to captionRefer to caption
(a) Case of IB dataset
Refer to captionRefer to caption
(b) Case of N1 dataset
Figure 5: Case studies on IB and N1 datasets. Three graphs are the origin, the coarsening, and the conversion, from left to right. The heatmap below illustrates the weight matrix of the classifier, where the x-axis denotes the feature dimensions, and the y-axis corresponds to three graphs from 0 to 5 (each takes up two rows). Best viewed in color.
We conduct case studies on IB and N1 datasets to cover each category and explore the choices of the classifier after MSLgo. The results are in Figure 5.
As shown in Figure 5a, high-weighted dimensions gather in the coarsening graph view, contributing more scores to the prediction. It reveals the case of prominent clique structures in social networks like IB. Meanwhile, we find the centrality of the conversion graph view in Figure 5b, which states the fitness of line graphs with edge-central modeling when meeting loops and chains in bioinformatics like N1.
In addition to unique structures, we can still find in the heatmap that the classifier retains weights for some feature dimensions in the original graph view. It is because, after the constrained coarsening and conversion, some nodes retained in the graph still hold valuable structural information, such as the bridge node connecting two structures and the edge node connecting inside and outside a clique. They should be paid equal attention to those unique components.
6 Conclusion
In this paper, we propose a novel multi-view graph structural representation learning model via graph coarsening (MSLgo) on GT architectures for graph classification task. We focus on loops and cliques and investigate the feasibility of treating particular structures node-like as a whole. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of MSLgo on six real-world datasets, and the experimental results demonstrate that MSLgo can outperform SOTAs in graph classification task.