### v2

# Node-like as a Whole: Structure-aware Searching and Coarsening for Graph Classification

Abstract.
Graph Transformers (GTs) have made remarkable achievements in graph-level tasks. However, most existing works regard graph structures as a form of guidance or bias for enhancing node representations, which focuses on node-central perspectives and lacks explicit representations of edges and structures. One natural question is, can we treat graph structures node-like as a whole to learn high-level features? Through experimental analysis, we explore the feasibility of this assumption. Based on our findings, we propose a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc) on GT architecture for graph classification. Specifically, we build three unique views, original, coarsening, and conversion, to learn a thorough structural representation. We compress loops and cliques via hierarchical heuristic graph coarsening and restrict them with well-designed constraints, which builds the coarsening view to learn high-level interactions between structures. We also introduce line graphs for edge embeddings and switch to edge-central perspective to construct the conversion view. Experiments on eight real-world datasets demonstrate the improvements of GRLsc over 28 baselines from various architectures.
Graph Classification, Graph Coarsening, Graph Transformer
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Theory of computation Graph algorithms analysisccs: Computing methodologies Supervised learning
1. Introduction
Graph Neural Networks (GNNs) have become a significant approach to Graph Representation Learning (GRL) recently, achieving remarkable performance on various node-level (such as node classification (Chen et al., 2023a) and link prediction (Chamberlain et al., 2023)) and graph-level (such as graph classification (Niepert et al., 2016; Xinyi and Chen, 2019; Yao et al., 2023)) tasks. GNNs model the graph structure implicitly, leveraging graph topology information to aggregate neighborhood node features. However, Xu et al. (2019) have proved that the aggregation mechanism of GNNs has local limitations, which has encouraged more researchers to move from GNNs to GTs (Liu et al., 2023a; Chen et al., 2023b; Wu et al., 2023). GTs utilize position encoding to represent the graph topology globally, jumping out of the neighborhood restriction and interacting between distant node pairs. Although GTs can learn long-distance structural information, most GNNs and GTs use graph structures as guidance (Huo et al., 2023) or bias (Ying et al., 2021) to obtain better node representations rather than directly representing them.
Structural information is crucial for graph-level tasks, especially Graph Structural Learning (GSL) (Zhang et al., 2023c; Zou et al., 2023; Zhang et al., 2023a). GSL aims to learn the optimized graph structure and representation from the original graph, further improving the model performance on downstream tasks (e.g. graph classification). Kernel-based methods (Shervashidze et al., 2009, 2011) lead the way of GSL in the early stage, which selects representative graph structures carefully to test the isomorphism and update weights in classifiers. However, this category of methods suffers due to computational bottlenecks of the pair-wise similarity (Wu et al., 2021; Yang et al., 2023). With the development of deep learning, researchers have proposed lots of new graph learning paradigms such as Graph Contrastive Learning (GCL) (Sun et al., 2020; Ma et al., 2023; Liu et al., 2023b). Using various data augmentation strategies like node dropping (Ma et al., 2023), edge perturbation (Zhang et al., 2023b), attribute masking (Yuan et al., 2023), and subgraph sampling (Liu et al., 2023b), GCL ensures the semantic matching between views and enhances the robustness of the model against multifaceted noises. Though GCL has made considerable achievements, views obtained through the strategies above keep the original structures, which restricts model capabilities since limited usage of high-level structural information.
To fully utilize topological information and capture high-level structures, one natural question arises as to whether we can treat graph structures node-like as a whole for graph-level tasks. Take the molecule graph as an example, specific structures such as molecular fragments and functional groups contain rich semantics. Random perturbation of these structures will produce additional structural information, which has been proved to be invalid nevertheless (Zhu et al., 2023). Some researchers have focused on this problem and defined the overall treatment idea as Graph Compression (Liu et al., 2022), Pooling (Wei et al., 2021; Baek et al., 2021; Lv et al., 2023), and Coarsening (Zhang et al., 2023b). Inspired by Convolutional Neural Networks (CNNs) to compress the set of vectors into a compact representation (Wu et al., 2022a), this series of works aims to address the flat architecture of GNNs, emphasizing the hierarchical structures of graphs (Ying et al., 2018).
Figure 1. Pre-experiment on a toy molecule graph set, which shows four graphs with different features, all containing a benzene ring. (B) illustrates the pairwise comparison between four graphs. (C) presents the comparison between the graph and its coarsening view. We leverage two measurements of Euclidean Distance (bar plots, lower for closer) and Cosine Similarity (line plots, higher for closer) to reveal the latent relations between graphs.
We further explore the feasibility of treating graph structures node-like as a whole through a set of pre-experiments. Figure 1 shows the results. The pre-experiments follow the plain message-passing mechanism and summation readout function. We train the pre-experimental model on four toy graphs and obtain representations of the whole graph, benzene ring structure, and main distinguishing atom sets. We measure the spatial distance and angular separation between representations by Euclidean Distance and Cosine Similarity, respectively. Figure 1(B) shows the comparison between different graphs. We divide the graph pairs into three groups according to different atom positions and different atom numbers. They are: (1) same position, different numbers: AC and BD; (2) different positions, same number: AB and CD; (3) different positions, different numbers: AD and BC. The experiments present instructive results, where the gap between loop is less, w.r.t. graph and diff. It highlights that the prime factor causing the differences lies behind the main distinguish atom sets rather than the benzene ring. Thus, we compress the benzene ring into one node with the same settings and obtain the coarsening view of each graph. Figure 1(C) shows the distance between the original graph and its coarsening view. Pairs such as A and C, whose main distinguish atom sets are arranged adjacent, leave a relatively closer cosine similarity, while B and D show a significant difference. It indicates that simple coarsening will lose part of structural information. In summary, pre-experiments tell us two enlightenments: (1) some structures contribute relatively less to distinguish graphs, which can further turn into a coarsening view to magnify the high-level structural information; (2) after treating structures node-like as a whole, some of the structural information may be traceless, calling for additional consideration for the relative position of neighborhoods.
Next, we reinforce the conclusions drawn from the pre-experiment with other real-world scenarios. Many systems, like recommendation (Yan et al., 2024), education (Rehm et al., 2024), and e-commerce (Fan et al., 2023), reveal imbalanced distributions of nodes, edges, and attributes, forming specific structures noteworthy. For example, cliques are groups of nodes with a high internal closeness and low external (Tsitsulin et al., 2023). From the node-central view, it is necessary to analyze the internal connections. Rich connections provide sufficient information for node-level analysis, such as Meanwhile, standing on the system side, the sparse edges outside cliques may generate high-level structural information, and many works (Xu et al., 2019; Ying et al., 2018) have proved that models can benefit from them. Low-density edges outside cliques represent how two cliques connect. It can answer some interesting questions, such as: what other sports will the group interested in football be interested in, which goods will the group who buy more electronic products buy, and so on.
Therefore, based on the above discovery, we focus on loops and cliques and introduce two views: a coarsening view and a conversion view to emphasize individual components of the graph. The former coarsens graph structures into one node based on clustering to learn high-level structural information. The latter transforms nodes and edges from the node-central perspective to edge-central, which highlights relative position and is complementary to the loss of coarsening. Specifically, we construct the graph coarsening view through a heuristic algorithm restricted by well-designed constraints and build the line graph conversion view to augment relative position information. Together with the original graph view, we train separate GT for each view for graph encoding. Finally, we concatenate multi-view representation vectors as the final embedding of an entire graph.
The contributions of this paper are as follows:
  • We design a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc), which leverages three views to learn more comprehensive structural information for downstream graph classification tasks.
  • We propose a hierarchical heuristic algorithm to compress the loops and cliques and construct the graph coarsening view, which captures high-level topological structures.
  • We introduce the line graph conversion view, which retains relative position information between neighbor nodes.
  • We verify the performance of GRLsc on 8 datasets from 3 categories, compared with 28 baselines from 6 categories. GRLsc achieves better results than SOTAs.
2. Related Work
2.1. Graph Pooling.
Graph Pooling (Wei et al., 2021; Baek et al., 2021; Lv et al., 2023) is one of the pivotal techniques for graph representation learning, aiming to compress the input graph into a smaller one. Due to the hierarchy characteristics of graphs, hierarchical pooling methods have become mainstream nowadays. Methods fall into two categories depending on whether breaking nodes. The first category conducts new nodes via graph convolution. The pioneer in hierarchical pooling, DIFFPOOL (Ying et al., 2018), aggregates nodes layer by layer to learn high-level representations. Inspired by the Computer Vision field, CGKS (Zhang et al., 2023b) constructs the pooling pyramid under the contrastive learning framework, extending the Laplacian Eignmaps with negative sampling. MPool (Islam et al., 2023) takes advantage of motifs to capture high-level structures based on motif adjacency. The second category condenses graphs by sampling from old nodes. SAGPool (Lee et al., 2019) masks nodes based on top-rank selection to obtain the attention mask subgraphs. CoGSL (Liu et al., 2022) extracts views by perturbation, reducing the mutual information for compression. MVPool (Zhang et al., 2023a) reranks nodes across different views with the collaboration of attention mechanism, then selects particular node sets to preserve underlying topological.
2.2. Graph Coarsening and Condensation.
Similar to Graph Pooling, both Graph Coarsening (Pang et al., 2021; Chen et al., 2023c; Zhang and Zhao, 2024) and Graph Condensation (Jin et al., 2022b, a) are graph reduction methods that simplify graphs while preserving essential characteristics (Hashemi et al., 2024). Generally speaking, Graph Coarsening groups and clusters nodes into super node sets using specified aggregation algorithms. It preserves specific graph properties considered high-level structural information, revealing hierarchical views of the original graph. L19 (Loukas, 2019) proposes a restricted spectral approximation with relative eigenvalue error. SS (Zhang and Zhao, 2024) conducts both intra- and inter-cluster features to strengthen the fractal structures discrimination power of models. In addition to the above spectral methods, researchers try to find other ways to make measurements. KGC (Chen et al., 2023c) calculates graphs equipped with Gromov-Wasserstein distance.
Graph Condensation, first introduced in (Jin et al., 2022b), leverages learnable downstream tasks’ information to minimize the loss between the original graph and some synthetic graphs. They propose a method based on gradient matching techniques, called GCond (Jin et al., 2022b), to condense the structures via MLP. However, GCond involves a nested loop in steps. Additional requirements for scalability led to the creation of DosCond (Jin et al., 2022a), EXGC (Fang et al., 2024), and many other works (Li et al., 2023; Zhang et al., 2024). DosCond (Jin et al., 2022a) is the first work focusing on graph classification tasks via graph condensation. EXGC (Fang et al., 2024), on the other hand, heuristically identifies two problems affecting graph condensation based on gradient matching and proposed solutions. Both types of work hope to achieve the same downstream task effect as the original graph while reducing the graph scale. For a further detailed understanding of those techniques, we recommend reading the survey (Hashemi et al., 2024).
2.3. Graph Transformers.
Graph Transformers (GTs) (Chen et al., 2023b; Wu et al., 2023) alleviate the problems of over-smoothing and local limitations of GNNs, which has attracted the attention of researchers. The self-attention mechanism learns long-distance interactions between every node pair and shows tremendous potential for various scenarios. Graphormer (Ying et al., 2021) integrates edges and structures into Transformers as biases, outperforming in graph-level prediction tasks. U2GNN (Nguyen et al., 2022) alternates aggregation function leveraging Transformer architecture. Exphormer (Shirzad et al., 2023) builds an expansive GT leveraging a sparse attention mechanism with virtual node generation and graph expansion. In addition, due to the quadratic computational constraint $O(N^{2})$ of self-attention, several works (Wu et al., 2022b; Liu et al., 2023a) have been proposed to focus on a scalable and efficient Transformer. The former puts forward a new propagation strategy adapting arbitrary nodes in scale. The latter combines pooling blocks before multi-head attention to shrink the size of fully connected layers. To summarize, we design our model based on the capabilities of GTs, which can effectively aggregate features of distant nodes using attention mechanisms and overcome the limitations of local neighborhoods.
Figure 2. Framework of our proposed model GRLsc (center). We show an example input graph $G_{i}$ and the corresponding intermediate output coarsening graph (bottom right). We will explain the coarsening and conversion details later.
3. Preliminaries
3.1. Notations
3.1.1. Graphs.
Given a graph $G=\{\mathcal{V},\mathcal{E}\}$ where $\mathcal{V}=\{v_{i}\}_{i=1}^{N}$ and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ denote the set of nodes and edges, respectively. We leverage $A\in\{0,1\}^{N\times N}$ to indicate the adjacency matrix, and $A[i,j]=1$ when there is an edge between node $v_{i}$ and $v_{j}$. We use $H=\{h_{ij}\}^{N\times D}$ to denote the node features, where $D$ shows the dimension of feature space, and $h_{ij}$ represents the feature of node $v_{i}$ in dimension $j$.
3.1.2. Problem Definition.
For supervised graph representation learning, given a set of graphs $\mathcal{G}=\{G_{1},\cdots,G_{M}\}$ and their labels $\mathcal{Y}=\{y_{1},\cdots,y_{M}\}$, our goal is to learn a representation vector $H_{i}$ for each graph, which can be used in downstream classification tasks to predict the label correctly.
3.2. Universal Graph Transformer (U2GNN)
U2GNN (Nguyen et al., 2022) is a GNN-based model that follows the essential aggregation and readout pooling functions. Xu et al. (2019) claim that a well-designed aggregation function can further improve the performance. Thus, U2GNN replaces the plain strategy which GNN uses with the following aggregation function:
(1) $h_{t,v}^{(k)}=\mathbf{AGG}_{\mathsf{att}}(h_{t-1,v}^{(k)}),$
where $h_{t,v}^{(k)}$ denotes the representation vector of node $v$ in step $t$ of layer $k$, which aggregates from step $t-1$. It provides a powerful method based on the transformer self-attention architecture to learn graph representation. We take U2GNN as the backbone and cover the details later in Section 4.4.
4. Methodology
4.1. Overview
Figure 2 shows the framework diagram of our proposed model GRLsc, which mainly contains three parts: multi-view construction of graphs, graph encoders, and classifier (omitted in the figure). Given an input graph $G_{i}$, GRLsc constructs three views step by step. Keep the original graph as $G_{o}$, and coarsen $G_{i}$ via the Loop and Clique Coarsening (LCC) Block to build view $G_{lcc}$. After that, we design the Line Graph Conversion (LGC) Block to transfer $G_{lcc}$ to view $G_{lgc}$. Subsequently, GRLsc trains a GT for each view separately to obtain graph-level representation, which is directly concatenated and fed into the downstream classifier.
This chapter is structured as follows: Section 4.2 and Section 4.3 describe LCC Block and LGC Block, respectively. Section 4.4 explains the rest of the details of GRLsc.
4.2. Loop and Clique Coarsening
4.2.1. Algorithm Description
Algorithm 1 Loop and Clique Coarsening (LCC) Block
0:  the input graph $G_{i}$, the max loop length $\delta$, the max hierarchy depth $\sigma$;
0:  the coarsening view $G_{lcc}$;
1:  $G_{lcc}\xleftarrow{}G_{i},G_{cc}\xleftarrow{}\emptyset,vis\xleftarrow{}\emptyset,V_{lcc}\xleftarrow{}V_{i}$;
2:  sort $V_{lcc}$ from high degrees to low;
3:  while not all nodes in $vis$ do
4:     $cur\xleftarrow{}V_{lcc}[0]$;
5:     fetch a clique $Cli$ not visited containing node $cur$ ;
6:     $G_{cc}\xleftarrow{update}Cli,vis\xleftarrow{}vis\cup Cli\setminus\{cur\}$;
7:     if exist node set $rest$ in the neighborhoods of $cur$ but has not been visited yet then
8:        find cliques $Cli_{rest}$ in $\sigma$ depth according to $rest$;
9:        $G_{cc}\xleftarrow{update}Cli_{rest},vis\xleftarrow{}vis\cup Cli_{rest}\setminus\{cur\}$;
10:     else
11:        $vis\xleftarrow{}vis\cup cur$;
12:     end if
13:  end while
14:  if less or no cliques found then
15:     $G_{lc}\xleftarrow{}\emptyset$;
16:     for all $v_{i}\in G_{lcc}$ do
17:        find all loops $L_{i}$ in $\delta$ length beginning with $v_{i}$;
18:        $G_{lc}\xleftarrow{update}L_{i}$;
19:     end for
20:     $G_{lcc}\xleftarrow{update}\{G_{cc},G_{lc}\}$;
21:  end if
22:  return $G_{lcc}$
From pre-experiments, we know that treating graph structures node-like as a whole via graph coarsening is possible. Existing methods mainly pay attention to the hierarchical structure of graphs: some cluster nodes based on graph convolution to do graph coarsening (Ying et al., 2018), and some implement node condensing by iterative sampling (Lee et al., 2019). Aggregation of first-order adjacent nodes has intuitive interpretability, which means connections between neighborhoods, while a deeper coarsening shows a poor explanation. Moreover, in high-level coarsening, the main distinguishing node sets disappear into clusters, which leads to an ambiguity for topology. Thus, GRLsc achieves graph coarsening of loops and cliques with shallow coarse-grained clustering restricted by two hard constraints, which magnifies high-level structural information while preserving the characteristic nodes in graphs to the greatest extent.
We first give formal definitions of loops and cliques.
Definition 4.1 (Loop).
Given an undirected graph $G$, a path $G_{l}=\{v_{1}v_{2}\cdots v_{n}v_{n+1}\}$ is a loop if there is no repetitive node in $\{v_{1},\cdots,v_{n},v_{n+1}\}$ except $v_{1}=v_{n+1}$ and $(v_{k},v_{k+1})\in\mathcal{E}_{G}$ for $k=1,2,\cdots,n$. $G_{l}$ is a k-loop if $|\mathcal{V}_{G_{l}}|=k$.
Definition 4.2 (Clique).
Given an undirected graph $G$, a subgraph $G_{c}$ is a clique if $\forall v_{i},v_{j}\in\mathcal{V}_{G_{c}},(v_{i},v_{j})\in\mathcal{E}_{G}(i\neq j)$. $G_{c}$ is a k-clique if $|\mathcal{V}_{G_{c}}|=k$.
Coarsening requires counting all loops and cliques in the graph. However, according to the definitions above, loops are contained within cliques. For example, given a 4-clique marked $Cli=\{ABCD\}$, we can find four 3-loops: $\{ABC\}$, $\{ABD\}$, $\{ACD\}$, and $\{BCD\}$, and one 4-loop $\{ABCD\}$, except for one 4-clique, which is not we want, since identifying $Cli$ as a clique is much more straightforward than five loops. Moreover, since the Maximum Clique Problem (MCP) is NP-hard (Bomze et al., 1999), finding all cliques is also NP-hard. Algorithms require an approximation or a shortcut due to enormous search space.
Algorithm 1 describes how the LCC Block works. Based on Tarjan (Trajan, 1974) and Clique Percolation (CP) (Palla et al., 2005), LCC heuristically iterates the graph hierarchy using loop length and hierarchy depth as constraints for pruning. LCC finds cliques under depths less than $\sigma$ (lines 3-13) first, counting loops only when finding fewer or no cliques (lines 14-21). When updating graphs, LCC reconstructs graphs to build the coarsening view. Mathematically, given an original graph $G_{o}=\{\mathcal{V},\mathcal{E}\}$, we aim to rebuild an intermediate coarsening graph $G_{lcc}$ with $n$ nodes where $n\ll|\mathcal{V}|$. In methods (Loukas and Vandergheynst, 2018; Loukas, 2019), supernode $v_{i}$ in $G_{lcc}$ aggregates nodes in $G_{o}$ according to node partitioning sets $\mathcal{P}=\{\mathcal{P}_{1},\cdots,\mathcal{P}_{n}\}$. We use indication matrix $M_{p}\in\{0,1\}^{n\times|\mathcal{V}|}$ to denote the aggregation based on partition set $\mathcal{P}$, where $m_{ij}=1$ if node $j$ in partition $\mathcal{P}_{i}$. We use a new adjacency matrix to represent $G_{lcc}$:
(2) $A_{lcc}=M_{p}AM_{p}^{T},$
where we multiply $M_{p}$ to the right to ensure the construction of the undirected graph. Here, we notice that $DEG=diag(A_{lcc})$ gives weight $a_{ii}$ for each supernode $i$, representing the sum of all node degrees corresponding to the partition $\mathcal{P}_{i}$. To directly extract and learn high-level structural information, we consider making a separation. We assign the weight of supernodes to 1 with $DEG^{-1}=diag(a_{ii}^{-1},\cdots,a_{nn}^{-1})$ instead of summation. The normalized adjacency matrix is
(3) $\bar{A}_{lcc}=diag(a_{ii}^{-1},\cdots,a_{nn}^{-1})(M_{p}AM_{p}^{T}),$
leaving each supernode equal contribution to the high-level structure initially. Similarly, we can define the node features of $G_{lcc}$ using $M_{p}$:
(4) $H_{lcc}=M_{p}H,$
where $H_{*}$ indicates the feature representation of *. This equation is equivalent to summing the nodes according to the partition sets.
So far, the focus of graph coarsening falls on how to obtain the partitioning set $\mathcal{P}$. LCC finds partitioning sets by searching for loops and cliques, where each $\mathcal{P}_{i}$ represents a loop or a clique. In other words, the process of LCC is to find a linear transformation matrix $A_{L}$ about $A$, such that
(5) $A^{\prime}=A_{L}AA_{L}^{T}=\begin{bmatrix}\mathcal{P}_{1}&E_{12}&\cdots&E_{1n}\\ E_{21}&\mathcal{P}_{2}&\ddots&\vdots\\ \vdots&\ddots&\ddots&\vdots\\ E_{n1}&\cdots&\cdots&\mathcal{P}_{n}\end{bmatrix},$
where the diagonal $\mathcal{P}_{i}=\{1\}^{k_{i}\times k_{i}}$ is the partitioning set containing $k_{i}$ nodes, and the non-diagonal $E_{ij}\in\{0,1\}^{k_{i}\times k_{j}}$ records the connection between partition $\mathcal{P}_{i}$ and partition $\mathcal{P}_{j}$. Next, by mapping each submatrix to 0 or 1 with the indicator function $\mathcal{I}(X)$, we can get the same result as Equation 2, i.e,
(6) $A_{lcc}=\mathcal{I}(A^{\prime}),$
(7) $\mathcal{I}(X)=\begin{cases}k_{i}&X=\mathcal{P}_{i}\\ 1&\exists x\in X,X=E_{ij},x=1\\ 0&\text{otherwise}\end{cases}.$
Figure 3. LCC Block contains two procedures: clique coarsening and loop coarsening. We design two constraints, hierarchy depth, and loop length, to control the graph coarsening process. A-D in the blue enclosure shows the details of the clique coarsening under hierarchy depth constraint, ①-③ in the red the loop coarsening with loop length constraint.
Figure 3 further depicts the workflow of the LCC Block. We design two coarsening algorithms with hierarchy depth constraint for clique and loop length constraint for loop. We go into details one by one.
4.2.2. Clique Coarsening with Hierarchy Depth Constraint.
The NP-hard nature of finding all cliques determines the impossibility of exhaustive search, especially for large graphs. GRLsc takes advantage of the hierarchical characteristic of graphs, taking nodes with the highest degree (Alg 1 line 2) hop by hop to find cliques. We set a distance $\sigma$ to control the depth of the recursion. A-D inside the blue closure in Figure 3 demonstrates the clique coarsening process with hierarchy depth constraints. (A) Given an example input graph with a clique structure, consider the node $M$ with the highest degree as the center node. (B) Search for possible cliques formed by connections within 1-hop neighbors of the central node and coarsen the found clique structure to rebuild a new node while preserving the central node. In step C, the central node $M$ remains unchanged, and the supernode $A=\{A_{1},A_{2},A_{3},M\}$. (C) Set the hierarchy depth constraint $\sigma$ and stop the coarsening of the current central node when the search range exceeds the limitation. If $\sigma=1$, the searching process simplifies to DFS. (D) If there exist nodes having 1-hop neighbors not searched yet (Alg 1 line 7), switch the central node and repeat the flow A-D until all nodes and edges are covered. For example, $C_{1},C_{2}$ are around node $B$, located outside the $\sigma$-hop of the central node $M$. After switching the central node to $B$, we can continue to search for the cliques.
4.2.3. Loop Coarsening with Loop Length Constraint.
Not all loops are suitable for coarsening. Long loops may reveal two chains or sequences interacting (e.g., backbone in proteins (Wang et al., 2023)). The semantics will change if coarsening happens. GRLsc sets a range $\delta$ for the maximum detection length (Alg  1 line 17). We default it to 6 to cover usual cases, such as squares and benzene rings. ①-③ inside the red closure in Figure 3 gives the loop coarsening process with loop length constraint. ① Given an example graph with loops of different lengths. ② Pick the starting node and transform the graph into a DFS sequence. ③ Set the loop length constraint $\delta$. When the sequence length exceeds the constraint, we can make the pruning on that chain. For example, the sample graph contains two 4-loop: $\{ABEF\}$ and $\{BCDE\}$ and one 6-loop: $\{ABCDEF\}$. When $\delta=4,5$, only $\{ABEF\}$ and $\{BCDE\}$ will be selected, forming a coarsening graph consisting of two supernodes. When $\delta=6$, $\{ABCDEF\}$ is added to the candidate set, replacing two loops above to construct one supernode. Although the triangle does not appear in the sample graph, it is a plain structure in the real world. We can handle it either as a 3-loop or a 3-clique.
4.2.4. Time Complexity
According to the previous description, LCC Block contains two parts, clique and loop coarsening, and the total time complexity of the algorithm is the sum of them. Given the input graph $G=\{\mathcal{V},\mathcal{E}\}$, we search for cliques following the hierarchy depth $\sigma$, marking nodes visited if no possible neighbors can form the new cliques. The time complexity of clique coarsening is $\mathcal{O}(\mathcal{V}+\mathcal{E})$. $\sigma$ only gives an order for searching and does not influence the linear computational cost. As for loop coarsening, we traverse along the DFS sequence and only keep loops under length constraints $\delta$. The time complexity of loop coarsening is $\mathcal{O}(\delta\mathcal{V}+\mathcal{E})$. For each node, we look back within $\delta$-level seeking loops satisfying the needs. Since $\delta$ is relatively small, the computation stays linear, $\mathcal{O}(\mathcal{V}+\mathcal{E})$. Therefore, the time complexity of the final LCC Block is $\mathcal{O}(2(\mathcal{V}+\mathcal{E}))$, achieving a linear growth in the scale of the input graph.
4.2.5. Limitation
Both loops and cliques are common structures in real-world graph networks. Understanding these structures is beneficial to thoroughly learning the network. LCC Block focuses on solving loops and cliques, so it is limited to other networks where these structures do not exist (e.g., long chains). To effectively alleviate this limitation of LCC, we introduce LGC Block next section 4.3 to further supplement the structural information through the conversion.
4.3. Line Graph Conversion
Figure 4. LGC Block workflow. Left side A-B shows the line graph conversion with hard coarsening examples (a.1 and a.2). Right side ①-③ illustrates how the input graphs with different position information are distinguished by LGC but not LCC.
This section tells why to build the LGC view. Line graph is a usual method for edge representations in node-level tasks such as POI recommendation (Chanpuriya et al., 2023). Edge augmentation is a considerable view to make a thorough representation integrating various architectures and scenarios (Mo and Yamana, 2023; Liu et al., 2023a). Therefore, we build LGC to attach an edge-central perspective in graph-level tasks. Firstly, the definition of line graphs is as follows.
Definition 4.3 (Line Graph).
Given an undirected graph $G_{o}=\{\mathcal{V},\mathcal{E}\}$, $L(G)=\{L(\mathcal{E}),\mathcal{E}_{L(G)}\}$ is the line graph of $G$, where $L(\mathcal{E})$ and $\mathcal{E}_{L(G)}$ denote the node set and edge set of the line graph, and $L$ is a mapping function from the edge set of the original graph to the node set of the line graph, satisfying $\forall e_{ij}\in\mathcal{E},L(e_{ij})\in L(\mathcal{E})$, and $\forall e_{ij},e_{ik}\in\mathcal{E},(L(e_{ij}),L(e_{ik}))\in\mathcal{E}_{L(G)}(j\neq k)$.
In plain words, a line graph turns the edges in the original graph into new vertices and the endpoints of two edges in the original graph into new edges. This explicit transformation can reduce the difficulty of relative position modeling, which is implicit in the node-central view of the original graph. We can build LGC through existing tools (Hagberg et al., 2008) conveniently. For those graphs with no edge attributes, we sum features of endpoints for consistency following Equation 4. In other words, we takes edges as the partitioning set and further aggregates features to build the LGC view. Mathematically, given an input graph $G_{o}$ and its feature vector $H$, LGC computes:
(8) $H_{lgc}=M_{e}[L(H_{lcc})]=M_{e}[L(M_{p}H)],$
where $L(\cdot)$ is the conversion function according to definition 4.3, $M_{e}$ is the indication matrix with edge partitions, and $H_{lcc}$ is the feature representations output by Equation 4.
In previous sections, we know LCC is limited if no loops and cliques exist, and the model suffers relative position information loss if we treat structures node-like. LGC is a supplementary of the coarsening view. We explain this from two aspects: some hard-coarsening examples and different position information of neighborhood nodes. Figure 4 shows the details.
4.3.1. LGC Block with Hard Coarsening Examples.
GRLsc only focuses on loops and cliques. However, not all graphs contain either structures, such as long chains. A simple copy view contributes less if we leave these hard-coarsening examples alone. LGC transforms these graphs into a fittable form that is easy for postprocessing while preserving the original structural information. For example, A-B in Figure 4 shows the LGC workflow with some hard coarsening examples. (A) Given a sample input graph without loops and cliques, we exclude two hard-coarsening structures: (a.1) Claw-like structure, consisting of a central node and three (or more) independent nodes connected to it. (a.2) Long chain with $n(=3)$ nodes connected one by one. Neither of them can be handled efficiently by LCC. (B) Convert structures through LGC Block to build an edge-centric view. In a claw-like structure, all edges are connected through the central node, forming a new clique after conversion. The long chain with $n$ nodes can decline the length to $n-1$ in one LGC calculation, reducing the scale effectively. GRLsc applies one LGC step on the intermediate coarsening graph of LCC, retaining the newly generated structures after conversion.
4.3.2. Different Position Information.
From pre-experiments, we know that relative position suffers if we condense graph structures into one node. Though plain coarsening strategies decrease the graph scale, it makes positions vague, such as between Graphs A and B in Figure 1. In essence, position information is the relative position among nodes connected by node-edge sequence. LGC models edge-central positional relationships, describing structural information by adding a new view to obtain a more comprehensive representation. We use ①-③ on the right of Figure 4 for illustration. ① Given two sample input graphs, each contains six nodes, four of which form a square, and the other two (denoted by A and B in the figure) connect to the square. The two sample graphs differ only in connection positions, one at adjacent nodes in the square and the other at diagonal nodes. ② If the square is identified as a 4-loop and coarsened to a supernode by LCC, it can not distinguish the two graphs by connection cases. ③ Realize the conversion to edge-central view by LGC, which can preserve different position information and differentiate graphs with subtle structural differences. As shown in Figure 4, LGC can identify the local claw-like structure containing node $B$ and convert it into a new triangle. We will get two LGC views with different positions of the triangle of node $B$ (in blue closure) adjacent or opposite to the triangle of $A$.
4.3.3. Limitation
Although LGC is complementary to LCC, it still has limitations. LGC can help LCC deal with graphs that don’t have loops and cliques, but line graph conversion is mostly a complication of the graph (increasing the scale). For a fully connected graph $G$ with $n$ nodes, the line graph $L(G)$ contains $\frac{n(n-1)}{2}$ nodes. It is an unacceptable growth rate at a quadratic level. LGC has poor adaptability to dense graphs, and its performance decreases when the number of edges in the intermediate coarsening graph by LCC far exceeds the number of nodes.
4.4. Our Model: GRLsc
After acquiring three views, we can build our model GRLsc based on U2GNN (Nguyen et al., 2022). As shown in Figure 2, GRLsc takes $G_{i}$ as input and obtains the coarsening view $G_{lcc}$ and line graph conversion view $G_{lgc}$ with LCC and LGC Block, respectively. For $G_{lcc}=\{\mathcal{V}_{lcc},\mathcal{E}_{lcc}\}$, GRLsc calculates node embeddings $h$ according to Equation 1 and Equation 4 and uses the summation readout function to obtain graph embedding $H_{lcc}$. Specifically, it is
(9) $H_{lcc}=\sum_{v\in\mathcal{V}_{lcc}}[h_{0,v}^{1};\cdots;h_{0,v}^{K}],$
where $h_{0,v}^{k}$ denotes the initial embedding of node $v$ in layer $k$. For each layer $k$, GRLsc iterates $T$ steps aggregating $N$ sampling neighbors and passes $h_{T,v}^{k}$ to the next layer $k+1$.
We use the same operation for the other two views, $G_{o}$ and $G_{lgc}$, and achieve corresponding embeddings, $H_{o}$ and $H_{lgc}$. We apply concatenation across views to obtain the final embedding $H_{i}$ for the input graph $G_{i}$ as
(10) $H_{i}=[H_{o};H_{lcc};H_{lgc}].$
After that, we feed the embedding $H_{i}$ to a single fully connected (FC) layer.
(11) $\hat{y}_{i}=WH_{i}+b,$
where $W$ is the weight matrix, and $b$ is the bias parameter, adapting to the embedding dimension increase. The loss function is cross-entropy as follows:
(12) $loss=-\sum_{i=1}^{M}y_{i}\log(\sigma(\hat{y}_{i})),$
where $\sigma$ denotes the softmax function. For further implementation details, we suggest to refer to  (Nguyen et al., 2022).
5. Experiments
5.1. Experimental Settings
5.1.1. Datasets.
Table 1. The statistics of datasets from TUDataset. Avg.N, Avg.E, and Avg.D denote the average number of nodes, edges, and degrees.
Dataset # Graphs # Classes Avg.N Avg.E Avg.D Category
CO 5000 3 74.49 2457.78 37.39 Social Network
IB 1000 2 19.77 96.53 8.89 Social Network
IM 1500 3 13.00 65.94 8.10 Social Network
DD 1178 2 284.32 715.66 4.98 Bioinformatics
N1 4110 2 29.87 32.30 2.16 Molecules
PTC 344 2 25.56 25.96 1.99 Bioinformatics
PRO 1113 2 39.06 72.82 3.73 Bioinformatics
N109 4127 2 29.68 32.13 2.16 Molecules
We evaluate our approach on eight widely used datasets from TUDataset (Morris et al., 2020), including three social network datasets: COLLAB (CO), IMDB-BINARY (IB), and IMDB-MULTI (IM), two molecules datasets: NCI1 (N1) and NCI109 (N109), and three bioinformatics datasets: D&D (DD), PTC_MR (PTC), and PROTEINS (PRO). Table 1 shows the details of the datasets.
  • COLLAB is a scientific collaboration dataset in physics. A graph corresponds to a researcher’s ego network, which has three possible labels indicating the fields that the research belongs to.
  • IMDB-BINARY (MULTI) shows a movie collaboration network between actors/actresses and movies in IMDB. An edge connects two people when they participate in the same scene. The labels of BINARY are Action and Romance, and MULTI’s are Comedy, Romance, and Sci-Fi.
  • NCI1 (109) is relative to anti-cancer screens. Each input graph represents a chemical compound, treating atoms and bonds as nodes and edges. Classification tasks reveal whether the chemicals appear positive or not. NCI1 aims for non-small cells and NCI109 for ovarian cells.
  • D&D is a dataset of proteins based on the same data as PROTEINS, where nodes represent secondary structure elements with their type information and edges show spatial proximity.
  • PTC_MR reports the carcinogenicity for rats of 344 chemical compounds.
  • PROTEINS is a dataset of proteins. In each graph, nodes represent the amino acids, and edges exist if two nodes are less than 6 Angstroms apart. It includes two categories: enzymes and non-enzymes.
Table 2. The main results of the graph classification task on eight datasets (mean accuracy (%) and standard deviation). The best scores are in bold and underline. ’-’ indicates that results are unavailable in the original or published papers. A.R shows the average ranking of each method. We highlight every method with outstanding performance in each category.
Method CO IB IM DD N1 PTC PRO N109 A.R
GW $72.84^{0.28}$ $65.87^{0.98}$ $43.89^{0.38}$ $78.45^{0.26}$ $62.30^{0.30}$ $57.26^{1.41}$ $71.67^{0.55}$ 18.71
WL $79.02^{1.77}$ $73.40^{4.63}$ $49.33^{4.75}$ $79.78^{0.36}$ $\underline{\bf{82.19}}^{0.18}$ $57.97^{0.49}$ $74.68^{0.49}$ $\underline{\bf{82.46}}^{0.24}$ 8.75
GCN $81.72^{1.64}$ $73.30^{5.29}$ $51.20^{5.13}$ $79.12^{3.07}$ $76.00^{0.90}$ $59.40^{10.30}$ $75.65^{3.24}$ $67.09^{3.43}$ 9.13
GAT $75.80^{1.60}$ $70.50^{2.30}$ $47.80^{3.10}$ $74.40^{0.30}$ $74.90^{0.10}$ $66.70^{5.10}$ $74.70^{2.20}$ 14.57
GraphSAGE $79.70^{1.70}$ $72.40^{3.60}$ $49.90^{5.00}$ $65.80^{4.90}$ $63.90^{7.70}$ $65.90^{2.70}$ $64.67^{2.41}$ 14.67
DGCNN $73.76^{0.49}$ $70.03^{0.86}$ $47.83^{0.85}$ $79.37^{0.94}$ $74.44^{0.47}$ $58.59^{2.47}$ $75.54^{0.94}$ 14.00
CapsGNN $79.62^{0.91}$ $73.10^{4.83}$ $50.27^{2.65}$ $75.38^{4.17}$ $78.35^{1.55}$ $66.00^{1.80}$ $76.28^{3.63}$ 9.57
GIN $80.20^{1.90}$ $75.10^{5.10}$ $52.30^{2.80}$ $75.20^{2.90}$ $79.10^{1.40}$ $64.60^{7.00}$ $76.20^{2.80}$ $68.44^{1.89}$ 7.75
GDGIN - - - $77.80^{3.60}$ $60.30^{4.50}$ $73.70^{3.40}$ 15.33
GraphMAE $80.32^{0.46}$ $75.52^{0.66}$ $51.63^{0.52}$ $80.40^{0.30}$ $75.30^{0.39}$ 6.00
InfoGraph $70.65^{1.13}$ $73.03^{0.87}$ $49.69^{0.53}$ $72.85^{1.78}$ $73.80^{0.70}$ $61.65^{1.43}$ $74.40^{0.30}$ 15.14
JOAO $75.53^{0.18}$ $70.21^{3.08}$ $47.22^{0.41}$ $75.81^{0.73}$ $74.86^{0.39}$ $73.31^{0.48}$ 17.50
RGCL $70.92^{0.65}$ $71.85^{0.84}$ $78.86^{0.48}$ $75.03^{0.43}$ 15.75
CGKS $76.80^{0.10}$ $79.10^{0.20}$ $63.50^{1.30}$ $76.00^{0.20}$ 7.75
DIFFPOOL $75.48^{0.00}$ $73.14^{0.70}$ $49.53^{3.98}$ $80.64^{0.00}$ $62.32^{1.90}$ $76.25^{0.00}$ $61.98^{1.98}$ 11.00
SAGPool $78.85^{0.56}$ $72.55^{1.28}$ $49.33^{4.90}$ $76.45^{0.97}$ $74.18^{1.20}$ $71.86^{0.97}$ $74.06^{0.78}$ 14.71
ASAP $78.64^{0.50}$ $72.81^{0.50}$ $50.78^{0.75}$ $76.87^{0.70}$ $71.48^{0.42}$ $74.19^{0.79}$ $70.07^{0.55}$ 13.43
GMT $80.74^{0.54}$ $73.48^{0.76}$ $50.66^{0.82}$ $78.72^{0.59}$ $75.09^{0.59}$ 9.60
SAEPool - - - - $74.48^{0.00}$ $80.36^{0.00}$ $75.85^{0.00}$ 6.67
MVPool - $52.00^{0.80}$ $82.20^{1.40}$ $80.10^{1.30}$ $\underline{\bf{85.70}}^{1.20}$ $81.90^{1.60}$ 6.20
PAS - - $53.13^{4.49}$ $79.62^{1.75}$ $77.36^{3.69}$ 5.33
GraphGPS - - - - $79.44^{0.65}$ $53.75^{6.20}$ $76.27^{0.95}$ 12.67
U2GNN $77.84^{1.48}$ $77.04^{3.45}$ $53.60^{3.53}$ $80.23^{1.48}$ $69.63^{3.60}$ $78.53^{4.07}$ 4.17
UGT - - - - $77.55^{0.16}$ $80.12^{0.32}$ $75.45^{1.26}$ 6.33
CGIPool $80.30^{0.69}$ $72.40^{0.87}$ $51.45^{0.65}$ $78.62^{1.04}$ $74.10^{2.31}$ $77.94^{1.37}$ 9.83
DosCond - - - $78.92^{0.64}$ $71.70^{0.20}$ 13.00
KGC - $69.20^{1.37}$ $61.58^{2.49}$ $66.43^{0.92}$ 18.33
SS - $73.89^{0.70}$ $51.80^{1.20}$ $79.78^{0.49}$ $76.13^{0.60}$ 6.00
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $79.08^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $76.37^{4.37}$ $76.98^{1.90}$ 2.63
5.1.2. Baselines.
To fully evaluate the effectiveness of our model, we select 28 related works from 6 categories for comparison.
As for the Graph Transformer framework, we pick 17 baselines. They are (I) 2 kernel-based methods: GK (Shervashidze et al., 2009) and WL (Shervashidze et al., 2011); (II) 8 GNN-based methods: GCN (Kipf and Welling, 2017), GAT (Velickovic et al., 2018), GraphSAGE (Hamilton et al., 2017), DGCNN (Zhang et al., 2018), CapsGNN (Xinyi and Chen, 2019), GIN (Xu et al., 2019), GDGIN (Kong et al., 2022), and GraphMAE (Hou et al., 2022); (III) 4 Contrastive Learning methods: InfoGraph (Sun et al., 2020), JOAO (You et al., 2021), RGCL (Shuai et al., 2022), and CGKS (Zhang et al., 2023b); and (IV) 3 GT-based methods: GraphGPS (Rampásek et al., 2022), U2GNN (Nguyen et al., 2022), and UGT(Hoang and Lee, 2024).
We also choose 11 models from 2 categories to verify our Graph Coarsening technique. One is (V) 7 Graph Pooling methods: DIFFPOOL (Ying et al., 2018), SAGPool(Lee et al., 2019), ASAP (Ranjan et al., 2020), GMT (Baek et al., 2021), SAEPool (Zhu et al., 2022), MVPool (Zhang et al., 2023a), and PAS (Wei et al., 2024). The other is (VI) 4 Graph Coarsening and Condensing methods: CGIPool (Pang et al., 2021), DosCond (Jin et al., 2022a), KGC (Chen et al., 2023c), and SS (Zhang and Zhao, 2024).
To further explain the unique design of our LCC block, we select 3 coarsening techniques: networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Each one takes two variants: neighborhood and clique. We will explain the details later in Section 5.4.
5.1.3. Implementation Details.
We follow the work (Xu et al., 2019; Nguyen et al., 2022) to evaluate the performance of our proposed method, adopting accuracy as the evaluation metric for graph classification tasks. To ensure a fair comparison between methods, we use the same data splits and 10-fold cross-validation accuracy to report the performance. In detail, we set the batch size to 4 and the dropout to 0.5. We pick the hierarchy depth in $\{1,2,3\}$, the number of Transformer Encoder layers in $\{1,2,3,4\}$, the number of sampling neighbors in $\{4,8,16\}$, the hidden size in $\{128,256,512,1024\}$, and the initial learning rate in $\{5e^{-4},1e^{-4},5e^{-4},1e^{-3}\}$. We utilize Adam (Kingma and Ba, 2015) as the optimizer. All experiments are trained and evaluated on an NVIDIA RTX 3050 OEM 8GB GPU and 16GB CPU. Our code is available: https://github.com/NickSkyyy/LCC-for-GC.
5.2. Results and Analysis
5.2.1. Main Results.
Table 2 shows the main results of the graph classification task. GRLsc outperforms other baseline models on most datasets, with the highest average ranking 2.63 above all. In detail, the boost of GRLsc is higher in social networks than in molecules and bioinformatics. WL, GraphMAE, CGKS, PAS, U2GNN, and SS show competitive results among their categories, even better than GRLsc (e.g., WL in N1 and N109). The main reason is that cliques of social networks are more distinct, while plain cliques and loop structures can not cover functional molecules and proteins completely. These main experimental results show the limitation of our model: the classification ability slightly drops when fewer loops and cliques appear. N1 and N109 contain many simple loops and long chains, while some unconnected graphs in PRO make it hard to coarsen and convert. We will show more details later in the ablation of the LCC block (Section 5.4) and case study (Section 5.6).
5.2.2. Robustness Studies.
We also investigate the robustness of GRLsc. Figure  5 presents the model performance of 10-fold in the best epoch for each dataset. Most experiment folds are higher than or equal to the average, having a mean turbulence of 2.58%. The worst two are PTC and PRO, with standard deviations of 5.22% and 4.37%, respectively. In PTC, the performance gap between the best fold (1 and 4, 82.86%) and the worst fold (5 and 9, 67.65%) is 15.21%, about 5 in 34 graphs. The best two are CO and DD, with standard deviations of 0.42% and 0.78%, respectively. In CO, the performance gap between the best fold (1, 98.80%) and the worst fold (8, 97.20%) is 1.60%, about 8 in 500 graphs.
5.3. Block Ablation Study
In this section, we will evaluate the effectiveness of each component of our model. We design two sets of ablation experiments, one for modules to analyze the effectiveness of GRLsc. And the other to discuss the impact of different coarsening strategies on downstream classification tasks.
[…]
Figure 5. 10 folds / best epoch on every dataset. The red dashed line plots the average accuracy.
Table 3. Block ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
GRLsc $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\bf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\bf{76.37}}^{4.37}$ $\underline{\bf{76.98}}^{1.90}$
- ${}_{w/o\,LGC}$ $97.96^{0.56}$ $87.60^{2.58}$ $61.47^{3.10}$ $99.32^{1.06}$ $77.42^{1.64}$ $69.50^{6.24}$ $75.02^{4.12}$ $74.90^{2.05}$
- ${}_{w/o\,LCC}$ $97.60^{0.81}$ $87.10^{2.91}$ $58.60^{3.74}$ $83.78^{4.63}$ $75.33^{2.12}$ $66.92^{8.38}$ $74.48^{3.42}$ $72.79^{2.14}$
- ${}_{w/o\,LGC+LCC}$ $96.50^{0.84}$ $84.10^{2.91}$ $57.87^{3.98}$ $83.02^{5.33}$ $72.53^{2.28}$ $63.09^{7.90}$ $68.01^{3.13}$ $69.20^{1.04}$
5.4. Coarsening Strategy Ablation
To better discuss the impact of different coarsening strategies on the model performance, we select four categories of graph coarsening schemes: random, networkx (Hagberg et al., 2008), KGC (Chen et al., 2023c), and L19 (Loukas, 2019). Under each category, there are two variants: neighbor and clique.
  • neighbor is an algorithm based on adjacent structures, which mainly aggregates information according to 1-hop adjacent nodes. There is no such algorithm in networkx, so we pick nx.cycle instead. It is similar to clique but based on cycle searches.
  • clique is an algorithm based on various clique searches, which aggregate node information according to clique partitions.
Table 4. Coarsening strategy ablation studies on all eight datasets.
Method CO IB IM DD N1 PTC PRO N109
Neighbor $98.12^{0.65}$ $89.10^{4.09}$ $62.80^{3.45}$ $98.64^{1.94}$ $77.79^{1.48}$ $70.67^{6.53}$ $75.20^{3.99}$ $76.40^{1.99}$
Random $97.94^{0.58}$ $89.10^{3.78}$ $61.70^{1.90}$ $99.32^{0.74}$ $77.79^{2.36}$ $71.81^{4.61}$ $72.51^{5.68}$ $74.10^{1.80}$
nx.cycle $98.02^{0.60}$ $87.00^{2.31}$ $62.47^{4.06}$ $86.16^{4.10}$ $74.84^{2.59}$ $75.92^{6.31}$ $74.30^{4.57}$ $73.81^{2.21}$
nx.clique $98.04^{0.58}$ $87.90^{2.62}$ $62.53^{3.41}$ $86.19^{5.62}$ $75.69^{1.92}$ $72.05^{6.67}$ $72.77^{6.16}$ $73.88^{1.81}$
KGC.neighbor $98.08^{0.58}$ $87.00^{1.95}$ $58.87^{4.03}$ $97.71^{2.57}$ $76.18^{2.12}$ $75.00^{2.66}$ $73.13^{3.70}$ $73.42^{1.36}$
KGC.clique $96.06^{0.77}$ $89.30^{2.72}$ $59.40^{3.66}$ $82.00^{7.04}$ $74.60^{1.91}$ $70.61^{8.86}$ $74.75^{5.05}$ $72.84^{1.72}$
L19.neighbor $96.92^{0.84}$ $87.10^{2.62}$ $61.70^{3.17}$ $96.44^{4.82}$ $75.94^{2.20}$ $74.06^{8.07}$ $74.21^{4.45}$ $73.81^{2.14}$
L19.clique $97.28^{0.66}$ $88.20^{2.79}$ $61.87^{3.55}$ $91.08^{3.00}$ $77.81^{1.48}$ $71.20^{6.10}$ $74.30^{3.65}$ $75.14^{1.86}$
LCC $\underline{\bf{98.18}}^{0.42}$ $\underline{\bf{89.40}}^{2.50}$ $\underline{\bf{63.20}}^{3.55}$ $\underline{\bf{99.58}}^{0.78}$ $\underline{\textbf{79.08}}^{1.92}$ $\underline{\bf{76.43}}^{5.22}$ $\underline{\textbf{76.37}}^{4.37}$ $\underline{\textbf{76.98}}^{1.90}$
In terms of implementation, we set $K=5$ and $r=0.5$, according to the paper (Chen et al., 2023c; Loukas, 2019). The algorithm strategy is greedy, and other parameters are consistent with the origin. To ensure alignment with GRLsc, the maximum coarsening level is set to 1, which means only coarsening the original graph without higher-level compressions. The random algorithm sets $R=K=5$, and each supernode randomly selects components from the sampling sets until all nodes are covered. Table 4 shows the model performance under different coarsening strategies, where LCC achieves the optimal results under all eight datasets. It indicates that the graph coarsening view obtained by LCC is more suitable for graph classification tasks.
5.4.1. Time Analysis
Figure 6. Runtimes of all coarsening algorithms. We make logarithmic processing due to different scales of datasets.
Table 5. Scale analysis of graph coarsening. Each group represents the size of the intermediate coarsening graph under different strategies, where $\bar{V}$ and $\bar{E}$ represent the average nodes and edges, and $r_{V}$ and $r_{E}$ represent the scaling ratio of the coarsening graph compared to the original graph. The negative number means reducing the scale from the original graph, and the positive means expanding.
Method CO IB IM DD N1 PTC PRO N109
Original $\bar{V}$ 74.49 19.77 13.00 284.32 39.06 25.56 29.87 29.68
$\bar{E}$ 2457.50 96.53 65.94 715.66 72.82 25.96 32.30 32.13
nx.cycle $\bar{V}$ 74.49 19.77 13.00 284.32 29.84 25.30 39.06 29.65
$r_{V}$ 0.00 0.00 0.00 0.00 -0.24 -0.01 +0.31 0.00
$\bar{E}$ 2474.63 101.15 67.94 722.44 42.09 47.85 77.02 41.98
$r_{E}$ +0.01 +0.05 +0.03 +0.01 -0.42 +0.84 +1.38 +0.31
nx.clique $\bar{V}$ 42.46 3.77 2.05 299.08 32.31 25.86 40.79 32.16
$r_{V}$0.43 -0.81 -0.84 +0.05 -0.17 +0.01 +0.37 +0.08
$\bar{E}$ 7082.97 11.96 2.81 2788.05 109.06 102.75 259.48 108.95
$r_{E}$ +1.88 -0.88 -0.96 +2.90 +0.50 +2.96 +7.03 +2.39
KGC.neighbor $\bar{V}$ 5.24 2.05 1.51 90.15 13.86 12.49 13.31 13.78
$r_{V}$0.93 -0.90 -0.88 -0.68 -0.65 -0.51 -0.55 -0.54
$\bar{E}$ 19.17 1.47 0.64 448.42 24.53 18.66 37.13 24.37
$r_{E}$0.99 -0.98 -0.99 -0.37 -0.66 -0.28 +0.15 -0.24
KGC.clique $\bar{V}$ 10.91 3.26 1.92 116.00 17.33 16.31 17.74 17.27
$r_{V}$0.85 -0.84 -0.85 -0.59 -0.56 -0.36 -0.41 -0.42
$\bar{E}$ 60.55 5.38 1.80 506.74 27.54 22.76 46.28 27.41
$r_{E}$0.98 -0.94 -0.97 -0.29 -0.62 -0.12 +0.43 -0.15
L19.neighbor $\bar{V}$ 46.30 12.46 9.83 142.45 15.27 13.67 19.88 15.17
$r_{V}$0.38 -0.37 -0.24 -0.50 -0.61 -0.47 -0.33 -0.49
$\bar{E}$ 1275.87 57.22 46.81 516.42 25.57 21.26 47.52 25.40
$r_{E}$0.48 -0.41 -0.29 -0.28 -0.65 -0.18 +0.47 -0.21
L19.clique $\bar{V}$ 46.12 12.42 9.76 142.45 17.35 16.00 20.25 17.28
$r_{V}$0.38 -0.37 -0.25 -0.50 -0.56 -0.37 -0.32 -0.42
$\bar{E}$ 1290.52 57.92 45.58 535.58 27.55 22.83 50.80 27.43
$r_{E}$0.47 -0.40 -0.31 -0.25 -0.62 -0.12 +0.57 -0.15
LCC $\bar{V}$ 12.52 3.45 2.01 155.72 18.88 20.49 25.85 18.73
$r_{V}$0.83 -0.83 -0.85 -0.45 -0.52 -0.20 -0.13 -0.37
$\bar{E}$ 29.38 3.07 1.22 306.33 18.97 20.07 40.54 18.85
$r_{E}$0.99 -0.97 -0.98 -0.57 -0.74 -0.23 +0.26 -0.41
We analyze the linear time complexity of LCC in section 4.2.4. Here, we compare it with other methods. Figure 6 shows the results.
The results show that the runtime of LCC is significantly lower than other methods. Except for CO and DD, the computation time is within 2s for all other datasets. Runtime for CO is 7.60s, while for DD is 37.58s. The average runtime of KGC.neighbor is 44.55s, and L19.clique is 54.98s. Both of them are higher than LCC.
5.4.2. Scale Analysis
Though we do not require LCC to achieve the optimal coarsening effect, we still analyze and compare the scale of the intermediate coarsening graph $G_{lcc}$. We experiment with three other categories of strategies and their variants in addition to random algorithms. Given a set of original input graphs $\mathcal{G}_{o}=\{G_{o}^{1},\cdots,G_{o}^{M}\}$, we calculate and collate the scale of $\mathcal{G}_{lcc}=\{G_{lcc}^{1},\cdots,G_{lcc}^{M}\}$. We show average node and edge numbers through $\bar{V}=\frac{1}{M}\sum_{i}\mathcal{V}_{G_{i}}$ and $\bar{E}=\frac{1}{M}\sum_{i}\mathcal{E}_{G_{i}}$, and analyze the scale changes leveraging formula $r_{V}=\frac{\bar{V}_{lcc}-\bar{V}_{o}}{\bar{V}_{o}}$ and $r_{E}=\frac{\bar{E}_{lcc}-\bar{E}_{o}}{\bar{E}_{o}}$. Table 5 shows the results.
(a) Loops: N109 (77)
(b) Cliques: IB (14)
(c) Plain: PTC (5)
Figure 7. Visualization of different coarsening strategies. For (a) loops, (b) cliques, and (c) plain graphs without loops and cliques, we find corresponding sample graphs from datasets (graph ID in parentheses). In each group, row 1 col 1 is the original input graph, row 2 col 1 is the coarsening result of GRLsc, and the other columns are two variants of the method in the same category. Different node colors represent different node classes. The graph layout does not contain semantics such as rotation and symmetry but only shows the connections.
In general, the LCC we designed can achieve a considerable coarsening result. The coarsening rate at the node level is about 52.2%. It is slightly higher at the edge level, 57.9% or so. In baselines, the best one is KGC.neighbor, with a 70.5% compression rate at the node level and 54.7% at the edge level. We do not expect LCC to achieve similar high rates. When the coarsening process is to the maximum extent, amounts of high-level structural information are lost.
Table 6. Optimal parameter settings.
Hyper-parameters CO IB IM DD N1 PTC PRO N109
$K$ 1 1 3 2 1 3 2 1
$T$ 2 2 1 2 2 4 1 1
$N$ 16 16 4 16 16 16 4 8
$|H|$ 1024 1024 1024 1024 1024 1024 1024 1024
$Lr$ 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001 0.0001
$|B|$ 4 4 4 4 4 4 4 4
Figure 8. Hyper-parameters Study. We evaluate four hyper-parameters: hierarchy depth $K$, the number of Transformer Encoder layers $T$, the number of sampling neighbors $N$, and the hidden size $|H|$.
As the number of nodes decreases linearly, edges have a faster decay. We cannot mine enough structural information in very simple hypergraphs. Thus, LCC achieves the maximum balance between coarsening nodes and preserving structures.
5.4.3. Visualization
We further analyze the coarsening comparison using visualization. Figure 7 shows the results.
Figures 7(a) and  7(b) visualize the loops and cliques that GRLsc focuses on. It appears that GRLsc can obtain the coarsening graph representation closest to the original graph. We retain the connections between supernodes to the greatest extent and mine the high-level structural information thoroughly based on the coarsening procedures. Other algorithms lack balance in coarsening and structural representation. Random algorithms break away from the original structural semantics and pursue to cover the whole graph by specifying the number of nodes. KGC presents a more concise coarsening pattern. However, it fails to express high-level structural information and has poor adaptability to loops. L19 relaxes the restriction, not limiting nodes located in multiple partitioning sets, but still cannot refine the structures.
In addition, we analyze for plain graphs that do not contain loops and cliques, as shown in Figure 7(c). GRLsc returns the same graph $G_{lcc}$ as the original input $G_{o}$ through LCC and makes a further conversion via LGC. In this kind of graph, KGC can effectively identify other structures as claw-like ones and realize their coarsening process.
5.5. Hyper-parameters Study
In this section, we explore the hyper-parameters of our model. We conduct experiments with different settings for 4 hyper-parameters: $K$, $T$, $N$, and hidden size on all eight datasets. Table 6 shows the optimal parameter settings, and Figure 8 shows the results of hyper-parameters studies. Generally, the optimal parameter settings are not the same for different datasets due to various data characteristics. For a single hyper-parameter, we can observe some patterns in results.
As for $K$ and $T$, a deeper and more complicated model improves model performances and enhances the ability to capture complex structural information. However, as the hierarchy extends continuously, overfitting leads to a decline in classification accuracy. The most striking example is the plot of CO and DD on $T$, where the performance of GRLsc decreases substantially when the number of layers changes from 3 to 4.
As for $N$ and $|H|$, in most cases, the larger the number of sampling neighbors and hidden size, the better the model performance. It indicates that the increase in the sampling and hidden size within a particular range can promote the model to learn high-level structures thoroughly and clearly. Among all datasets, IM, PRO, and N109 do not conform to the rules summarized above. IM and PRO can capture long-distance structural information with the help of deeper hierarchy depth without sampling more neighbors. N109 achieves a higher neighbor coverage with only a small amount of sampling for its multi-loop distribution (see Section 5.6 below for more details).
5.6. Case Study
(a) CO (5)
(b) DD (0)
(c) IB (5)
(d) IM (268)
(e) N1 (45)
(f) N109 (540)
(g) PTC (7)
(h) PRO (6)
Figure 9. Case study for all eight datasets. Each case contains three graphs from left to right: the original graph view, the graph coarsening view, and the line graph conversion view. The heat plot below them is the classifier weights, with the x-axis representing the feature dimension and the y-axis representing the three graphs (each graph view takes rows of the number of graph classes).
We conduct case studies on all eight datasets to fully cover each category and explore the choices of the classifier after GRLsc. The results are in Figure 9.
It can be seen from the cases intuitively that different datasets have different structural characteristics. Even datasets in the same category also have differences in representation. For example, (a) CO, (c) IB, and (d) IM have strong clique structure distribution, graphs in (h) PRO contain multiple connected branches and the similar datasets (e) N1 and (f) N109 consist of multi-chain and multi-loop structure, respectively, and so on. GRLsc focuses on loops and cliques and thus performs well on datasets rich in such structures. As for loops, we take (f) N109 as an example, which contains three benzene rings connected in turn. LCC can identify such loop structures and coarsen them into three supernodes. The remaining two independent carbon atoms are then converted by LGC into a triangle, as shown in the figure. We can see from the heatmap that both LCC and LGC contain vital dimensions contributing most to the classification weights, where LCC in (35, 3), LGC in (6, 4), (24, 5), etc.
As for cliques, we take (c) IB as an example. It forms 12 independent cliques centered on the actor or actress represented by the purple node. A clique represents a collection of actors or actresses from one scene, and some actors or actresses may appear in more than one scene. LCC weakens the concept of nodes but highlights the clique structure by coarsening, strengthening the connection between scenes with corporate actors or actresses. LGC further emphasizes the structural information of the original graph. A clique under the edge-centric view shows a new representation of the purple center node. Compared to the previous case, for datasets with prominent clique structures, LCC contributes more to the weight than LGC, where LCC in (2, 2), (21, 3), (63, 3), and LGC in (13, 5), (46, 5), and so on.
In addition to examples with remarkable features, the dataset also contains (b) DD with a more complex structure and (e) N1 without loops and cliques. We can observe the unique role of LCC and LGC from them. The former coarsens the loops and cliques via LCC and significantly reduces the graph scale, while the latter shortens long chains and marks the position of special forks, such as claw-like structures, through LGC.
Across all eight datasets, we notice that (h) PRO has more unique properties. The original graph is possibly not connected, containing some independent connected branches or components. LCC and LGC start from one particular node, traversing the node and edge set completely. Therefore, multiple connected branches do not affect the operation of our algorithm.
Finally, we can still find in the heatmap that the classifier retains weights for some feature dimensions in the original graph view. It is because, after the constrained coarsening and conversion, some nodes retained in the graph still hold valuable structural information, such as the bridge node connecting two structures and the edge node connecting inside and outside a clique. They should be paid equal attention to those unique components.
6. Conclusion and Future Work
6.1. Conclusion
In this paper, we propose a novel multi-view graph representation learning model via structure-aware searching and coarsening (GRLsc) on GT architectures for graph classification task. We focus on loops and cliques and investigate the feasibility of treating particular structures node-like as a whole. We build three unique views via graph coarsening and line graph conversion, which helps to learn high-level structural information and strengthen relative position information. We evaluate the performance of GRLsc on six real-world datasets, and the experimental results demonstrate that GRLsc can outperform SOTAs in graph classification task.
6.2. Future Work
Though GRLsc achieves remarkable results, we still have a long way to go. In the future, there are two main directions for discussion. First, GRLsc prefers to deal with static graphs at present, while graphs in the real world are constantly changing over time. We try to introduce dynamic graphs. Second, GRLsc focuses on two specific structures: loops and cliques. However, graph structures are more complex and diverse. We will consider extending to general structures to mine richer information at a high level.